Historically, 5% of people who see this ad the first time click on it. What is the probability that this ad will receive 10 clicks after being shown to 50 new people? 100? 300?
1
Expert's answer
2020-02-24T11:07:01-0500

This is a binomial distribution with p=0.05, q=1-p ; q=1-0.05=0.95 and x=10

P(x)= (nCx)*p^x*q^n-x

i) probability when n=50

=(50C10)*0.05^10*0.95^40

=0.0001289

ii)probability when n= 100

=(100C10)*0.05^10*0.95^90

=0.0167

iii) probability when n =300

=(300C10)*0.05^10*0.95^290

=0.0473